Given the linear function f(x)=-1/2x -4
a) Find f(-4). My answer is -9/2.
b) Determine the coordinates of the x-intercept. My answer is x=-1/2.
c) Determine the coordinates of the y-intercept. My answer is no y intercept.
d) Explain the difference between the x-intercept and the y-intercept. Not sure as to this answer.
d) Determine the slope. My answer is m=0, b=-9/2.
e) Determine the slope of a line perpendicular to this line. Not sure
If you can help me as to if it's right and explain the ones I did not do, and how you solved it.
Thanks

Answer by Alan3354(69443)

Given the linear function f(x)=-1/2x -4
Assuming you mean (-1/2)x - 4:
a) Find f(-4). My answer is -9/2.
f(-4) = (-1/2)*-4 - 4 = +2 - 4 = -2
-------------------
b) Determine the coordinates of the x-intercept. My answer is x=-1/2.
@y = 0: -x/2 - 4 = 0
x = -8 or (-8,0)
-------------
c) Determine the coordinates of the y-intercept. My answer is no y intercept.
@ x = 0: y = -4
(-4,0)
------------
d) Explain the difference between the x-intercept and the y-intercept. Not sure as to this answer.
At the x-int, y=0 and vice versa.
---------------------
d) Determine the slope. My answer is m=0, b=-9/2.
m = the coefficient of x = -1/2. b wasn't asked for, but it's -4.
--------------------
e) Determine the slope of a line perpendicular to this line. Not sure
The slope of lines perpendicular is the negative inverse of -1/2
m = +2